OpenStudy (anonymous):

Given the equation P = 2w + 2l solve for w.

OpenStudy (anonymous):

\[w = \frac{ P }{ 2 }- l\]

OpenStudy (anonymous):

2w = P - 2l B) w = 2l + P 2 C) w = P -2l 2 D) w = 2l - P 2 Eliminate